# Data-Copilot: Bridging Billions of Data and Humans with Autonomous Workflow

### v4

@@ 2.2 Interface Dispatch
To accurately comprehend user requests, Data-Copilot first parses the time, location, data object, and output format of user requests, which are critical to data-related tasks. For example, if the question is: "I want to compare the GDP and CPI trend in our area over the past five years", Data-Copilot parses it as: "Draw a line chart of China’s national GDP and CPI per quarter from May 2017 to May 2023 for comparison". To achieve this, we first invoke an external API to obtain the local time and network IP address, then feed this external information into LLM along with the original request to generate the parsed result.
Planing Workflow
Once the user’s intent is accurately understood, Data-Copilot plans a reasonable workflow to process the user’s request. We feed the user request and related interface descriptions into LLMs for the automatic deployment of workflows. Which interfaces should be invoked, and in what sequence, is entirely determined by Data-Copilot, based on the user’s request and the interface.
To achieve this, Data-Copilot generates a response in JSON format representing each step of the scheduling: ###Interface Call: {step1={"arg":"", "function":"", "output":"" ,"description":""}, step2={..}, ..}. Please refer to Section F.2 for a detailed prompt. Data-Copilot meticulously orchestrates the scheduling of the interface in either a sequential or parallel manner.
For instance, the first request in Figure 2: "Compare the return of CSI 300, GEM Index and CSI 1000 Index this year", Data-Copilot first plans five steps in its workflow. In the first two steps, it dispatches {Data Acquisition}, {Calcuate Return} in parallel for three indices. The last three steps successively invoke the {Visualization} to sequentially plot three indices on the same canvas. In the second case, Data-Copilot deploys a loop workflow to sequentially calculate the indicators of 50 constituent stocks by calling the {Loop and Rank}.
@@
Upon the deployment and execution of the workflow, Data-Copilot yields the desired results in the form of graphics, tables, and descriptive text. Additionally, it also provides a comprehensive summary of the entire workflow. This systematic summary not only provides clear and concise results but also sheds light on the steps taken to achieve those results, thereby enhancing transparency and understanding of user requests.
As the example shown in Figure G2, the request is "Forecasting China’s GDP growth rate...". Data-Copilot first interprets the user’s intent based on local time. Then, it deploys a three-step workflow: 1) Invoking {get-GDP-data} interface to acquire historical GDP data. 2) Invoking {predict-next-value} interface for forecasting. 3) Visualizing the output.
3 Qualitative Evaluation
3.1 Environment and Data Sources
We evaluate Data-Copilot utilizing Chinese financial market data, encompassing massive stocks, funds, economic data, real-time news, and company financial data. To closely mimic real-world scenes, we adopt the data interfaces (APIs) provided by Tushare [^1] https://tushare.pro/ to access data, which are commonly used by many financial companies. The vast and diverse array of financial time-series data enables a comprehensive evaluation of Data-Copilot’s performance. In the first phase, we use the gpt-4-0613222https://platform.openai.com/ for interface design and gpt-3.5-turbo-0613 for interface dispatch.
3.2 Seeds and Datasets
Human-proposed Request as Seed We invite 10 students in economics to submit 50 requests each. These requests should satisfy two requirements: 1) cover a wide range of common needs, including stocks, funds, and corporate. 2) involves different levels of difficulty. Then we filter out highly similar requests by computing the sentence embedding of each request. Lastly, we retain 173 high-quality requests, which are used as the seeds for Interface Design phase of the Data-Copilot.
Self-Request and Testing Benchmark As mentioned in Section 2.1, we employed a self-request strategy to expand the request set. Specifically, we feed 173 manually proposed seed requests along with all data descriptions and schema into GPT-4. We instruct GPT-4 to generate massive requests, covering diverse data. In this phase, we generate a total of 6480 requests. Then we compute the semantic similarity between all requests and filter out those with high similarity. Lastly, we retain 3547 requests. We randomly select 3000 requests from the total set for subsequent interface design. Meanwhile, the remaining 547 requests are treated as a test set. We manually annotate the final answers, task types, and workflow types for each testing instance. The statistics of three dataset are shown in Tables 1 and C1.
3.3 Qualitative Results
Data-Copilot designs many versatile interface tools through a self-request. As shown in Figure G3, each interface is clearly defined and versatile, including data acquisition, processing, dataframe manipulation, and visualization. These diverse interfaces endow the LLM with the capability to address complex requests. For instance, get-stock-prices-data interface integrates the functions of acquiring data in daily, weekly, and monthly frequencies, controlled by argument: freq, which is very similar to the process of programmer development.
Data-Copilot deploys workflows with different structures. Data-Copilot invokes the most appropriate interface and forms an efficient workflow with different structures, e.g., parallel, serial, or loop workflows, tailored to user requests. As shown in Figures G4 and G5, it employs a loop workflow for the requests: "The year-on-year net profit growth rate of the SSE50 ..." by invoking Loop-Rank interface. It iteratively calculates the financial indicators for each stock. Please refer to Appendix E for more visualization cases.
4 Quantitative Evaluation
@@ 4.1 Experiments Settings
Baselines 1. Global Planning (GP) (Shen et al., 2023; Qin et al., 2023a): We instruct the LLM to directly generate solving codes for user requests. 2. ReAct (Yao et al., 2022): The LLM resolves the user request using neural language and code step-by-step. 3. Reflexion (Shinn et al., 2023): LLM can also reflect on the feedback from the compiler at each step, progressively fulfilling user requests. For the Reflexion strategy, we limit it to a maximum of two rounds of reflection. 4. Multi-agent collaboration (Wu et al., 2023b; Hong et al., 2023; Liang et al., 2023a): We adopt three LLM-based agents, i.e., two coders and a manager, handling user requests simultaneously and outputting the final solution after mutual discussion. All methods are evaluated on the 547 testing instances (Section 3.2) three times. A detailed prompts of baselines is shown in Section F.3.
@@
Table 2: All methods use gpt-3.5 as based LLM.
@@
GP 50.8 28.5 823.6
ReAct 63.4 44.1 1515.9
Reflexion 75.7 59.0 2463.5
Multi-Agent 73.6 57.4 2835.2
Data-Copilot 79.9 70.2 561.2
@@ 4.2 Comparison Results
Data-Copilot significantly reduces the risk of failure. Compared to the best baseline (Reflexion), our method increases the success rate from 59% to 70.2%. In most cases, our system merely needs to invoke pre-designed interface tools and generate corresponding arguments to address user requests, significantly reducing the risk of errors. In contrast, baseline strategies must construct all code from scratch, which makes it easy to overlook certain details or make mistakes. Besides, the difference between Pass@1 and Exec@1 also indicates that baseline strategies often miss certain steps, whereas Data-Copilot can fully interpret user requests and methodically complete tasks step by step.
Data-Copilot reduces repetitive overhead for common requests. The consumed tokens represent computational overhead. Data-Copilot only consumes 561 tokens, which is just 30% of the consumption compared to baselines. Specifically, the output of Data-Copilot is very concise, only containing interface names and arguments. The primary token consumption lies in the input prompt which contains related tool descriptions. In contrast, baseline strategies may consume a substantial amount of tokens since they need to output a complete solving code for each request. In the real world, numerous requests are similar or even repetitive. Data-Copilot can reduce repetitive overhead by abstracting common demands into interface tools. However, baseline methods are compelled to repetitively generate code for each request.
@@ 4.3 Analysis
Data-Copilot exhibits a pronounced advantage on more challenging tasks. To better analyze the problem-solving capabilities of Data-Copilot, we categorize our dataset into three subsets based on the complexity of the workflow: Sequential, Parallel, and Complex. As shown in Table 3, for sequential tasks, the baselines and our method perform similarly, with only a 4% improvement. However, for the more challenging parallel or complex tasks, our improvements are substantial, with enhancements of 5.3% and 19%, respectively. We find that baseline methods often struggle to fully implement complex functionalities and may omit some critical steps, such as loop processing. In contrast, our methods effectively handle these tasks by invoking versatile interfaces and deploying workflow. These findings elucidate the advantages of Data-Copilot, which reduces task complexity.
We can deploy Data-Copilot locally using small LLMs with commendable performance. We analyze the performance of Data-Copilot with different LLMs. As shown in Figure 4, we test GPT family APIs and various open-source models. We observe GPT-4 surpasses all LLMs, achieving 75.6% Pass@1, while many open-source LLMs also achieve commendable results. For instance, codellama-13B achieves a 50.8% Pass@1, almost similar to the best baseline in Table 3 (Reflexion with gpt-3.5-turbo). This indicates that Data-Copilot can be deployed locally at a relatively low cost while maintaining commendable performance.
@@ 4.4 Human Evaluation
To assess whether the solutions align with human preferences, we perform human evaluations across two dimensions: rationality and conciseness. Specifically, we collect the solutions of Data-Copilot and Reflexion. Then we invite experts to compare which solution is better in terms of rationality and conciseness (detail in Appendix C). We present results in Figure 4. It demonstrates that our solution significantly outperforms the competitor in terms of rationality. Specifically, Data-Copilot prevails in 47.4% of instances, trailing behind the competitor in only 23.5%. Besides, as for conciseness, Data-Copilot lags slightly behind its competitor, with 27.8% wins against 30% losses. This is attributed to the fact that these pre-designed interfaces invoked by Data-Copilot are quite universal, accommodating a wide array of functions, which makes our system less concise. The results of human evaluation are consistent with the previous assessment. Although Data-Copilot sacrifices a bit of succinctness, it can generate more rational and correct solutions.
Figure 3: Results of Data-Copilot with various LLMs.
@@ 5 Conclusion
We propose a universal framework, Data-Copilot, to address a wide array of data-related tasks. It acts as a bridge between numerous heterogeneous data and humans, effectively managing, processing, and presenting data according to human tastes, reducing the dependence on tedious labor and expert knowledge. Acting like an experienced expert, Data-Copilot autonomously designs universal interface tools for various types of data and potential user demands and invokes these interfaces in real-time response for more reliable problem-solving.
6 Acknowledgement
This work is supported by the National Natural Science Foundation of China (No. 62376245), the Key Research and Development Program of Zhejiang Province, China (No. 2024C01034), the project of the Donghai Laboratory (Grant no. DH-2022ZY0013), and MOE Engineering Research Center of Digital Library.
@@ 6 Acknowledgement
Appendix A Discussion
A.1 What is Data-Copilot?
  • Data-Copilot is an efficient LLM-based data management agent, that processes massive data and rapidly responds to real-time human requests.
  • It contains interface design and invocation phases. Interface design involves a self-exploration phase for data sources to create many versatile interfaces.
  • For interface invocation, it parses human real-time requests and invokes the designed interfaces to retrieve and process relevant data and generates multi-form outputs, e.g., text, charts, and tables.
A.2 Why Data-Copilot is designed in two stages?
  • Reduces redundant generation: People often have a large number of similar or repetitive query needs, e.g., many people are interested in seeing which stock has the largest gain today. For these repetitive query requests, employing the LLM to write code from scratch for each user request would be highly inefficient and a significant waste of computational resources. To avoid inefficient and repetitive code generation for numerous similar requests, we abstract common demands into generic interfaces through data self-exploration in advance (Phase-1). For real-time requests (Phase-2), it only needs to invoke these interfaces, reducing redundant generation.
  • Performs better and consumes fewer tokens: Compared to a complete code, it only needs to invoke the relevant interfaces, which is simpler and less prone to errors. Our results show that it performs better and consumes fewer tokens.
  • Stronger interpretability: Compared to complex code, interface workflow is easier for humans to read and inspect.
Appendix B Experiments Details
During the interface design phase, we instruct LLM to generate Python code for each interface in the interface implementation. We filter out interfaces that cannot run properly at each iteration. For the interface Dispatch phase, we carry out hierarchical planning to improve efficiency: upon receiving a request, Data-Copilot first determines the type of data task involved (stock task, fund task, etc.) and then loads the corresponding interface description, followed by interface planning using the corresponding interface. We set the temperature to 0.2 for more consistent results.
Appendix C Human Evaluation Details
@@
Figure C1: We count the keywords for each request type (Stock, Corp., Fund, Others) in the test set.
@@ Appendix D Related Works
In the recent past, breakthroughs in large language models (LLMs) such as GPT-3, GPT-4, PaLM, and LLaMa (Brown et al., 2020; Chowdhery et al., 2022; Zhang et al., 2022; Zeng et al., 2023; Touvron et al., 2023; Ouyang et al., 2022; OpenAI, 2023b; Wei et al., 2022b; Wu et al., 2023c) have revolutionized the field of natural language processing (NLP). These models have showcased remarkable competencies in handling zero-shot and few-shot tasks along with complex tasks like mathematical and commonsense reasoning. The impressive capabilities of these LLMs can be attributed to their extensive training corpus, intensive computation, and alignment mechanism (Ouyang et al., 2022; Wang et al., 2022b; a).
@@
LLM-based Agent Recent studies have begun to explore the synergy between external tools and large language models (LLMs). Tool-enhanced studies  (Schick et al., 2023; Gao et al., 2022a; Qin et al., 2023a; Hao et al., 2023; Qin et al., 2023b) integrate external tools into LLM, thus augmenting the capability of LLMs to employ external tools. Several researchers have extended the scope of LLMs to include the other modality (Wu et al., 2023a; Surís et al., 2023; Shen et al., 2023; Liang et al., 2023b; Huang et al., 2023). In addition, there are many LLM-based agent applications (Xie et al., 2023), such as CAMEL (Li et al., 2023a), AutoGPT [^3] https://github.com/Significant-Gravitas/Auto-GPT, AgentGPT [^4] https://github.com/reworkd/AgentGPT, BabyAGI [^5] https://github.com/yoheinakajima/babyagi, BMTools [^6] https://github.com/OpenBMB/BMTools, LangChain [^7] https://github.com/hwchase17/langchain, Agentverse (Chen et al., 2023b), Autoagent (Chen et al., 2023a), MetaGPT (Hong et al., 2023), AutoGEN (Wu et al., 2023b), etc. Most of them are focused on daily tools or code generation and do not consider the specificity of data-related tasks. Except for learning to operate the tools, several contemporaneous studies (Cai et al., 2023; Qian et al., 2023) have proposed to empower LLMs to create new tools for specific scenarios like mathematical solving and reasoning. These impressive studies have revealed the great potential of LLM to handle specialized domain tasks.
@@
Applying LLM To Data Science Apart from these studies, the application of large models in the field of data science has garnered significant interest among researchers. FLAME (Joshi et al., 2023) investigates the feasibility of using NLP methods to manipulate Excel sheets. StructGPT (Jiang et al., 2023) explore reasoning abilities of LLM over structured data. LiDA (Dibia, 2023) and GPT4-Analyst (Cheng et al., 2023) focus on automated data exploration. Besides, many reseraches (Liu et al., 2023; Chang & Fosler-Lussier, 2023; Dong et al., 2023), like Sheet-Copilot (Li et al., 2023b), BIRD (Li et al., 2023c), DAIL-SQL (Gao et al., 2023), DIN-SQL (Pourreza & Rafiei, 2023), DB-Copilot (Wang et al., 2023b), MAC-SQL (Wang et al., 2023a), ChatDB (Hu et al., 2023), SQL-PaLM (Sun et al., 2023), and DB-GPT (Xue et al., 2023) apply LLMs to Text2SQL and table rasoning. Chain-of-Table (Wang et al., 2024) proposes a step-by-step reasoning strategy based on the table.
@@
(3) Distinct Implementation Approach: Data-Copilot designs its tools for data processing, rather than relying on SQL or pre-existing tools. This approach is more flexible and can meet complex user requirements.

### v5

@@ 2.2 Workflow Deployment
To accurately comprehend user requests, Data-Copilot first parses the time, location, data object, and output format of user requests, which are critical to data-related tasks. For example, the request is: "I want to compare the GDP and CPI trend in our area over the past five years", Data-Copilot parses it as: "Draw a line chart of China’s national GDP and CPI per quarter from May 2019 to May 2024 for comparison". To achieve this, we first invoke an external API to obtain the local time and network IP address, then feed this supportive information into LLMs along with the original request to generate the parsed result.
Planning Workflow
After providing the user request and a description of the relevant interfaces, Data-Copilot plans a logical interface workflow to process the user request. Which interfaces should be invoked, and in what sequence, is entirely determined by Data-Copilot. The LLMs are instructed to output in JSON format for each step invocation: Interface Call: {step1={"arg":"", "Interface":"", "output":"" ,"desc":""}, step2={..}, ..}.
For different requests, Data-Copilot designs interface workflows with different structures, such as sequential, parallel, or loop. For instance, the first request in Figure 2: "Compare the return of CSI 300, GEM Index and CSI 1000 Index this year", Data-Copilot plans a five-step workflow. In the first two steps, it dispatches Data Acquisition(), Calcuate Return() in parallel for three indices. The last three steps successively invoke the Visualization() to sequentially plot three indices. In the second case, it deploys a loop workflow by calling the Loop().
@@
Upon execution of the workflow, Data-Copilot yields the desired results in the form of graphics, tables, and descriptive text. Additionally, it also provides a comprehensive summary of the entire workflow. It greatly enhances the interpretability of the whole process, as the interface workflow is easy for humans to read and inspect. As the example shown in Figure G2, the request is "Forecasting China’s GDP growth rate...". Data-Copilot first interprets the user’s intent. Then it deploys a three-step workflow: 1) Invoking get-GDP-data() interface to acquire historical GDP data. 2) Invoking predict-next-value() interface for forecasting. 3) Visualizing the output.
@@
3 Qualitative Evaluation
3.1 Environment and Data Sources
We evaluate Data-Copilot utilizing Chinese financial market data, encompassing massive stocks, funds, economic data, real-time news, and company financial data. To closely mimic real-world scenes, we adopt the data interfaces (APIs) provided by Tushare [^1] https://tushare.pro/ to access data, which are commonly used by many financial companies. The vast and diverse array of financial time-series data enables a comprehensive evaluation of Data-Copilot’s performance. In the first phase, we use the gpt-4-0613 for data exploration design and gpt-3.5-turbo-0613 for workflow deployment.
3.2 Seeds and Datasets
Human-proposed Request as Seed We invite 10 students in economics to submit 50 requests each. These requests should satisfy two requirements: 1) cover a wide range of common needs, including stocks, funds, and corporate. 2) involves different levels of difficulty. Then we filter out highly similar requests by computing the sentence embedding of each request. Lastly, we retain 173 high-quality requests, which are used as the seeds for Data Exploration phase of the Data-Copilot.
en Self-request and Testing Benchmark As mentioned in Section 2.1, we employed a self-request to expand the request set. Specifically, we feed 173 human-proposed requests along with all data descriptions and schema into GPT-4. We instruct GPT-4 to generate massive requests. In this phase, we generate 6480 requests, then compute the semantic similarity between all requests, and filter out those with high similarity, and retain 3547 requests. We randomly select 3000 requests from the total set for subsequent interface design. Meanwhile, the remaining 547 requests are treated as a test set. We manually annotate the final answers, task types, and workflow types for each testing instance. The statistics of three datasets are shown in Tables 1 and B1.
3.3 Qualitative Results
Data-Copilot designs versatile interfaces through data exploration. In Figure G3, each interface is clearly defined and versatile, including data acquisition, processing, manipulation, and visualization. These diverse interfaces endow the LLM with the capability to address complex requests. For instance, get-stock-prices-data() interface integrates the functions of acquiring data by day, week, and month, controlled by argument: freq, which is similar to the process of programmer development.
Data-Copilot deploys workflows with different structures. Data-Copilot invokes the most appropriate interface and forms an efficient workflow with different structures, e.g., parallel, serial, or loop workflows, tailored to user requests. As shown in Figure 2, it employs a loop workflow for the requests: "The year-on-year net profit growth rate of the SSE50 ..." by invoking Loop-Rank interface. It iteratively calculates the financial indicators for each stock. Please refer to Appendix E for more visualization cases.
4 Quantitative Evaluation
@@ 4.1 Experiments Settings
Baselines 1. Vanilla LLM (Shen et al., 2023; Qin et al., 2023a): We instruct the LLM to directly generate solving codes for user requests. 2. ReAct (Yao et al., 2022): The LLM resolves the user request using neural language and code step-by-step. 3. Reflexion (Shinn et al., 2023): LLM can also reflect on the feedback from the compiler at each step, progressively fulfilling user requests. For the Reflexion strategy, we limit it to a maximum of two rounds of reflection. 4. Multi-agent collaboration (Wu et al., 2023b; Hong et al., 2023; Liang et al., 2023a): We adopt three LLM-based agents, i.e., two coders and a manager, handling user requests simultaneously and outputting the final solution after mutual discussion. All methods are evaluated on the 547 testing instances (Section 3.2) three times. The detailed prompts of all baselines are shown in Section F.3.
@@
Table 2: We demonstrate the accuracy and efficiency of all methods with gpt-3.5.
@@
[…]
@@ 4.2 Comparison Results
Data-Copilot significantly reduces the risk of failure. Compared to the best baseline (Reflexion), our method increases the success rate from 59% to 70.2%. In most cases, it merely needs to invoke pre-designed interface modules and generate corresponding arguments to address user requests, notably reducing the risk of errors. In contrast, baseline strategies must construct all code from scratch, making it easy to overlook certain details or make mistakes. Besides, the difference between Pass@1 and Exec@1 indicates that baseline strategies often miss certain steps, whereas Data-Copilot can fully interpret user requests and methodically complete tasks step by step.
Data-Copilot reduces repetitive overhead for common requests. The consumed tokens represent computational overhead. Data-Copilot only consumes 561 tokens, just 30% of the baseline’s consumption. Specifically, the output of Data-Copilot is very concise, only containing interface names and arguments. The primary token consumption lies in the input prompt which contains related interfaces descriptions. In contrast, baseline strategies may consume a lot of tokens since they need to output a complete solving code for each request. In the real world, numerous requests are similar or even repetitive. Data-Copilot can reduce repetitive overhead by abstracting common demands into interfaces. However, baseline methods have to repetitively generate code for each request.
@@ 4.3 Analysis
Data-Copilot exhibits a pronounced advantage on more challenging tasks. To better analyze the problem-solving capabilities of Data-Copilot, we categorize our dataset into three subsets based on the complexity of the workflow: Sequential, Parallel, and Complex. Complex workflows mean some intricate cases, mainly involving loops or nested workflows. As shown in Table 3, for sequential tasks, the baselines and our method perform similarly, with only a 4% improvement. However, for the more challenging parallel or complex tasks, our improvements are substantial, with enhancements of 5.3% and 19%, respectively. We find that baseline methods often struggle to fully implement complex functionalities and may omit some critical steps, such as loop processing. In contrast, our methods can effectively handle these tasks by invoking versatile interfaces and deploying workflow. These findings elucidate the advantages of Data-Copilot, which reduces task complexity.
We can deploy Data-Copilot locally using small LLMs with commendable performance. We analyze the performance of Data-Copilot with different LLMs. In Figure 4, we test GPT family and various open-source models. We observe GPT-4 surpasses all LLMs, achieving 75.6% Pass@1, while many open-source LLMs also achieve commendable results. For instance, codellama-13B achieves a 50.8% Pass@1, almost similar to the best baseline in Table 3 (Reflexion with gpt-3.5-turbo). This indicates that Data-Copilot can be deployed locally at a relatively low cost while maintaining commendable performance.
@@ 4.4 Human Evaluation
To assess whether the solutions align with human preferences, we perform human evaluations across two dimensions: rationality and conciseness. Specifically, we collect the solutions of Data-Copilot and Reflexion. Then we invite experts to compare which solution is better in terms of rationality and conciseness (detail in Appendix B). We present results in Figure 4. It demonstrates that our solution significantly outperforms the competitor in terms of rationality. Specifically, Data-Copilot prevails in 47.4% of instances, trailing behind the competitor in only 23.5%. Besides, as for conciseness, Data-Copilot lags behind its competitor, with 27.8% wins against 30% losses. This is attributed to the fact that these pre-designed interfaces invoked by Data-Copilot are universal, accommodating a wide array of functions, which makes our system less concise. The results of human evaluation are consistent with the previous assessment. Although Data-Copilot sacrifices a bit of succinctness, it can generate more rational and correct solutions.
4.5 Performance of other programming languages
Data-Copilot exhibits excellent scalability, allowing for easy switching to other programming languages by simply regenerating the corresponding interfaces. We tested three programming languages: Python, C++, and Matlab. The results indicate that Python performed the best (Pass@1: 70.2), followed by C++ (54.2), with Matlab (36.5) yielding the poorest results. Upon examination, we found that Matlab code often suffers from formatting errors or dimensional discrepancies in the data, rendering the program non-executable. We speculate this may be related to a lack of sufficient Matlab code in the pre-training corpus. So ultimately, we opted for Python for Data-Copilot.
Figure 3: Results of Data-Copilot using various LLMs as based model.
@@ 5 Conclusion
We propose a code-centric agent, Data-Copilot, to address a wide array of data-related tasks. Data-Copilot is a code-centric agent, which generates codes to process large-scale data, and generates generic interface modules in advance through a data exploration process to improve the success rate of real-time requests. It autonomously designs universal interface modules for various types of data and potential user demands and invokes these interfaces in real-time response for more reliable problem-solving. Our experiments show that Data-Copilot flexibly handles challenging requests about Chinese financial data, with a higher success rate and less token consumption.
@@
Appendix A Experiments Details
During the data exploration phase, we instruct LLM to generate Python code for each interface. Then we abstract interfaces from these requests. To merge the similar interfaces for each iteration, we adopt the following strategy:
1. Firstly, we calculate the semantic similarity between the names of all existing interfaces and the newly designed interface, then sort them by similarity from highest to lowest.
2. Then, we feed the top-10 existing interfaces into the LLM, and instruct the LLM to sequentially check whether they are functionally similar to the new interface, or can be merged using new parameters.
3. Lastly, we output the merged interface. If the LLM deems no merging necessary, the new interface is retained.
Besides, for the workflow deployment phase, we carry out hierarchical planning to improve efficiency: upon receiving a request, Data-Copilot first determines the type of data task involved (stock task, fund task, etc.) and then loads the corresponding interface description, followed by interface planning using the corresponding interface. We set the temperature to 0.2 for more consistent results.
Appendix B Human Evaluation Details
@@
Figure B1: We count the keywords for each request type (Stock, Corp., Fund, Others) in the test set.
Appendix C Limitations
Data-Copilot proposes a new paradigm for addressing the data-related task, through LLM. But we want to highlight that it still remains some limitations or improvement spaces:
1) Online Design Interface. The essence of Data-Copilot lies in effective interface design, a process that directly affects the effectiveness of subsequent interface deployments. Currently, this interface design process is conducted offline. Therefore, it is crucial to explore how to design the interface online and deploy it simultaneously. It greatly broadens the application scenarios of Data-Copilot.
2) System stability The interface deployment process can occasionally be unstable. The main source of this instability is because LLM is not fully controllable. Despite their proficiency in generating the text, LLMs occasionally fail to follow the instructions or provide incorrect answers, thus causing anomalies in the interface workflow. Consequently, finding methods to minimize these uncertainties during the interface dispatch process should be a key consideration in the future.
3) Data-Copilot possesses the potential to handle data from other domains effectively. Currently, our focus is on developing an LLM-based agent for the data domain. Due to the limited access to data, we chose the China Financial Data, which includes stocks, futures, finance, macroeconomics, and financial news. Although these data all belong to the financial domain, the data volume is extremely large, and the data schemas are highly different. The corresponding user’s requests are also diverse, which poses a great challenge to the current LLM. However, Data-Copilot has adeptly accomplished this task. Therefore, we believe Data-Copilot also possesses the potential to effectively handle data from other domains.
@@ Appendix D Related Works
In the recent past, breakthroughs in large language models (LLMs) such as GPT-3, GPT-4, PaLM, and LLaMa [Brown et al., 2020, Chowdhery et al., 2022, Zhang et al., 2022, Zeng et al., 2023, Touvron et al., 2023, Ouyang et al., 2022, OpenAI, 2023, Wei et al., 2022a] have revolutionized the field of natural language processing (NLP). These models have showcased remarkable competencies in handling zero-shot and few-shot tasks along with complex tasks like mathematical and commonsense reasoning. The impressive capabilities of these LLMs can be attributed to their extensive training corpus, intensive computation, and alignment mechanism [Ouyang et al., 2022, Wang et al., 2022b, a].
@@
LLM-based Agent Recent studies have begun to explore the synergy between external tools and large language models (LLMs). Tool-enhanced studies  [Schick et al., 2023, Gao et al., 2022, Qin et al., 2023a, Hao et al., 2023, Qin et al., 2023b] integrate external tools into LLM, thus augmenting the capability of LLMs to employ external tools. Several researchers have extended the scope of LLMs to include the other modality [Wu et al., 2023a, Surís et al., 2023, Shen et al., 2023, Liang et al., 2023b, Huang et al., 2023]. In addition, there are many LLM-based agent applications [Xie et al., 2023], such as CAMEL [Li et al., 2023a], AutoGPT [^2] https://github.com/Significant-Gravitas/Auto-GPT, AgentGPT [^3] https://github.com/reworkd/AgentGPT, BabyAGI [^4] https://github.com/yoheinakajima/babyagi, BMTools [^5] https://github.com/OpenBMB/BMTools, LangChain [^6] https://github.com/hwchase17/langchain, Agentverse [Chen et al., 2023b], Autoagent [Chen et al., 2023a], MetaGPT [Hong et al., 2023], AutoGEN [Wu et al., 2023b], etc. Most of them are focused on daily tools or code generation and do not consider the specificity of data-related tasks. Except for learning to operate the tools, several contemporaneous studies [Cai et al., 2023, Qian et al., 2023] have proposed to empower LLMs to create new tools for specific scenarios like mathematical solving and reasoning. These impressive studies have revealed the great potential of LLM to handle specialized domain tasks.
@@
Applying LLM To Data Science Apart from these studies, the application of large models in the field of data science has garnered significant interest among researchers. FLAME [Joshi et al., 2023] investigates the feasibility of using NLP methods to manipulate Excel sheets. StructGPT [Jiang et al., 2023] explore reasoning abilities of LLM over structured data. LiDA [Dibia, 2023] and GPT4-Analyst [Cheng et al., 2023] focus on automated data exploration. Besides, many reseraches [Liu et al., 2023, Chang and Fosler-Lussier, 2023, Dong et al., 2023], like Sheet-Copilot [Li et al., 2023b], BIRD [Li et al., 2023c], DAIL-SQL [Gao et al., 2023], DIN-SQL [Pourreza and Rafiei, 2023], PET-SQL [Li et al., 2024b], DB-Copilot [Wang et al., 2023b], MAC-SQL [Wang et al., 2023a], ACT-SQL [Zhang et al., 2023], ChatBI [Lian et al., 2024], CodeS [Li et al., 2024a], SQLPrompt [Sun et al., 2023a], ChatDB [Hu et al., 2023], SQL-PaLM [Sun et al., 2023b], and DB-GPT [Xue et al., 2023], EHRAgent [Shi et al., 2024] apply LLMs to Text2SQL and table rasoning. Chain-of-Table [Wang et al., 2024] proposes a step-by-step reasoning strategy based on the table.
@@
(3) Distinct Implementation Approach: Data-Copilot is a code-centric agent for data processing, rather than relying on SQL or pre-existing tools. This approach is more flexible and can meet complex user requirements.